The walls and ceiling inside your bedroom need to be painted. To save money, you decide that you will paint the bedroom yourself. Use the following information to solve this problem:
• The bedroom is 17 feet long by 18 feet wide, and the ceiling is 9 feet high.
• The inside of the bedroom door will be painted the same color as the walls.
• Two coats of paint will be applied to all of the painted surfaces.
• The room has one window, measuring 3 feet, 9 inches by 4 feet, which will not be painted.

Answer by Boreal(15235)

The surface area is
2 walls 18 x 9=162*2=324
2 walls 17 x 9=153*2=306
ceiling 17 x 18=306*1=306
That is 936 ft^2
subtract window or 4* 3 3/4=15
This is 921 ft^2
Two coats means the equivalent surface area as if the room were twice that or 1842 ft^2.
That is as far as one can go with the information.